# Well-posedness and stability of the Lagrange representation of the n-D wave equation via boundary triples

### v1

@@ 4. Stability
  • $\Gamma_{0}\neq\partial\Omega$ and that there exists $\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}$ open and non-empty such that $k_{2}>0$ on $\Gamma$ (i.e., there is an open set where the wave equation is damped).
@@
Making use of the results of the previous section, we have to show stability properties of the semigroup generated by $A_{\Theta}$. The notion of stabiliy we have in mind is the following one:
@@
A strongly continuous semigroup $(T(t))_{t\geq 0}$ with generator $G$ is called semi-uniformly stable, if and only if there exists a continuous non-increasing function $f\colon[0,\infty\rparen\to[0,\infty\rparen$ ) satisfying $\lim_{t\to\infty}f(t)=0$ such that for every $x\in\operatorname{dom}G$:
$\lim_{t\to\infty}\lVert T(t)x\rVert_{H}\leq f(t)\lVert x\rVert_{\operatorname{dom}G}$
@@
The notion of semi-uniform stability is nested in between strong stability where
@@
for all $x\in H$ is demanded and uniform stability, where
@@
We rely on the following criterium from [2, Thm. 1.1] [^13] The decay rate can be made explicit, cf. [6, Thm. 3.4] and its proof for details..
@@
Let $(T(t))_{t\geq 0}$ be a bounded strongly continuous semigroup with generator $G$ satisfying $\upsigma(G)\cap\mathrm{i}\mathbb{R}=\emptyset$. Then $(T(t))_{t\geq 0}$ is semi-uniformly stable.
@@
The embedding of $\operatorname{dom}A$ into $X$ is compact, i.e., $\operatorname{dom}A\overset{\mathsf{cpt}}{\hookrightarrow}X$.
@@
$\left\lVert\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{\operatorname{dom}A}^{2}\coloneqq\left\lVert\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{X}^{2}+\left\lVert A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{X}^{2}\leq C.$
@@
$\big{\lVert}\tfrac{1}{\rho}y_{n}\big{\rVert}_{\mathrm{L}^{2}(\Omega)}^{2}+\big{\lVert}T^{\frac{1}{2}}\nabla x_{n}\big{\rVert}_{\mathrm{L}^{2}(\Omega)}^{2}+\big{\lVert}k_{1}^{\frac{1}{2}}\gamma_{0}x_{n}\big{\rVert}_{\mathrm{L}^{2}(\tilde{\Gamma})}^{2}\\ +\big{\lVert}\tfrac{1}{\rho}\operatorname{div}T\nabla x_{n}\big{\rVert}^{2}_{\mathrm{L}^{2}(\Omega)}+\big{\lVert}T^{\frac{1}{2}}\nabla\tfrac{1}{\rho}y_{n}\big{\rVert}^{2}_{\mathrm{L}^{2}(\Omega)}+\big{\lVert}k_{1}^{\frac{1}{2}}\gamma_{0}\tfrac{1}{\rho}y_{n}\big{\rVert}^{2}_{\mathrm{L}^{2}(\tilde{\Gamma})}\leq C\text{.}$
@@
  • Since $(\tfrac{1}{\rho}y_{n})_{n\in\mathbb{N}}$ is bounded in $\mathrm{H}^{1}(\Omega)$, there exists a subsequence that converges in $\mathrm{L}^{2}(\Omega)$. Hence, $(\rho\tfrac{1}{\rho}y_{n(k)})_{k\in\mathbb{N}}$ converges in $\mathrm{L}^{2}(\Omega)$ as well.
  • Again since $(x_{n})_{n\in\mathbb{N}}$ is bounded in $\mathrm{H}^{1}(\Omega)$, there exists a subsequence that converges to $x$ in $\mathrm{L}^{2}(\Omega)$. W.l.o.g. we pass to that subsequence. To avoid the introduction of several (irrelevant) multiplicative constants, we use the symbol $\lesssim$ which stands for inequality up to a multiplicative constant independent of $n$. Note that appealing to the boundary condition we have
    $\lVert\gamma_{\nu}Tx_{n}\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}=\lVert k_{1}\gamma_{0}x_{n}+k_{2}\gamma_{0}\tfrac{1}{\rho}y_{n}\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}\lesssim C\text{.}$
    Integration by parts and the Cauchy–Schwarz inequality give
@@
    =$\stackrel{{\scriptstyle\mathllap{\text{Int.\ by parts}}}}{{=}}\big{\langle}-\operatorname{div}T\nabla(x_{n}-x_{m}),x_{n}-x_{m}\big{\rangle}+\langle\gamma_{\nu}T(x_{n}-x_{m}),\gamma_{0}(x_{n}-x_{m})\rangle$
    $\stackrel{{\scriptstyle\mathllap{\text{C.--S.\ ineq.}}}}{{\lesssim}}C\lVert x_{n}-x_{m}\rVert_{\mathrm{L}^{2}(\Omega)}+C\lVert\gamma_{0}(x_{n}-x_{m})\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})},$
    which shows that $(\nabla x_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $\mathrm{L}^{2}(\Omega)$ and hence convergent. By the closedness of the operator $\nabla$, we conclude that $x$ is also the limit of $(x_{n})_{n\in\mathbb{N}}$ in $\mathrm{H}^{1}(\Omega)$. Finally, closedness of $\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$ in $\mathrm{H}^{1}(\Omega)$ implies $x\in\mathring{\mathrm{H}}_{\Gamma_{0}}^{1}(\Omega)$. ∎
@@
The right hand side is of the form $1+K$, where $K$ is a compact operator. It is a consequence of the theorem of Riesz-Schauder, cf. [24, Thm. 6.2.1], that an operator $1+K$ is injective if and only if it is surjective. Now note that the left hand side is injective, since $\lambda-A_{\Theta}$ is injective and the resolvent is bijective. Thus $\lambda-A_{\Theta}$ is surjective. The open mapping theorem assures that $\lambda-A_{\Theta}$ is boundedly invertible. ∎
@@
If $\lambda\in\mathrm{i}\mathbb{R}$ is an eigenvalue of $A_{\Theta}$, the corresponding eigenvectors $x$ satisfy
$\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1}=0\quad\text{and}\quad k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2}=0.$
@@
We can calculate for $\lambda\in\upsigma(A_{\Theta})$ and $x\in\operatorname{dom}A_{\Theta}$ a corresponding eigenvector
@@
$=\tfrac{1}{2}\left(\langle Ax,x\rangle+\langle x,Ax\rangle\right)-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\tfrac{1}{2}\left(\langle B_{1}x,B_{2}x\rangle_{\mp}+\langle B_{2}x,B_{1}x\rangle_{\pm}\right)-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\operatorname{Re}\langle B_{1}x,B_{2}x\rangle_{\mp}-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\operatorname{Re}\langle\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp}-\operatorname{Re}\lambda\lVert x\rVert^{2}.$
If $\lambda\in\mathrm{i}\mathbb{R}$, then we obtain
$0=\operatorname{Re}\langle\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp}.$
Appealing to the boundary condition and the fact that the duality can be written as an inner product if the arguments are in $\mathrm{L}^{2}$ we infer
$0=\operatorname{Re}\langle k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp}=\operatorname{Re}\langle k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle=\lVert k_{2}^{\frac{1}{2}}\gamma_{0}\tfrac{1}{\rho}x_{2}\rVert^{2}\text{,}$
which implies $k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2}=0$. Employing the boundary condition again we deduce $\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1}=0$. ∎
Note that the previous result says that $\gamma_{0}\tfrac{1}{\rho}x_{2}=0$ on $\operatorname{supp}k_{2}$, which by assumption (A6) implies that there is a non-empty open set $\Gamma$ contained in $\operatorname{supp}k_{2}$ such that $\gamma_{0}\tfrac{1}{\rho}x_{2}=0$ on $\Gamma$.
@@
$\mathrm{i}\lambda u_{1}-\tfrac{1}{\rho}u_{2}&=0\quad\text{in}\;\mathrm{H}^{1}(\Omega)\\ \mathrm{i}\lambda u_{2}-\operatorname{div}T\nabla u_{1}&=0\quad\text{in}\;\mathrm{L}^{2}(\Omega).\end{aligned}$ (10)
@@
Let $\lambda\in\mathbb{R}\setminus\{0\}$ and let us assume that $\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}$ is an eigenvector to the eigenvalue $\lambda$. Then from Equation 10 we obtain that $u_{1},u_{2}\in\mathrm{H}^{1}(\Omega)$ and plugging the second equation into the first we obtain that $u_{1}\in\mathrm{H}^{1}(\Omega)$ has to satisfy
$\lambda^{2}\rho u_{1}+\operatorname{div}T\nabla u_{1}=0$
Combining the first equation of (10) and Lemma 4.6 gives
@@
Then the unique continuation principle [^14] It is here that our stronger assumptions from the beginning of the section come into play. They are not required anywhere else!, cf. Appendix A, implies that $u_{1}=0$ is the only solution, which in turn implies $u_{2}=0$ and therefore $u=0$, which is a contradiction. ∎
@@
Since $\operatorname{dom}A_{\Theta}\overset{\mathrm{cpt}}{\hookrightarrow}X$ by Proposition 4.4 it suffices to show that 00 is not an eigenvalue. Suppose 00 is an eigenvalue and $x\in\operatorname{dom}A_{\Theta}$ a corresponding eigenvector. Then from Equation 10 we immediately see $x_{2}=0$. Moreover, by Lemma 4.6
@@
$0=\langle\operatorname{div}T\nabla x_{1},x_{1}\rangle=-\langle T\nabla x_{1},\nabla x_{1}\rangle+\langle\gamma_{\nu}T\nabla x_{1},\gamma_{0}x_{1}\rangle.$
@@
Thus $x_{1}=0$ as well, but $x=0$ is not an eigenvector. Hence, 00 cannot be an eigenvalue. ∎
@@
By Theorem 3.7 $A_{\Theta}$ generates a contraction semigroup. Theorem 4.5 shows that $\sigma(A_{\Theta})$ consists only of eigenvalues, Lemma 4.7 and Lemma 4.8 show that the imaginary axis is contained in the resolvent set of $A_{\Theta}$. Proposition 4.3 then implies the claim. ∎
@@
We can also obtain semi-uniform stability of Equation 1 under suitable assumptions on the perturbation
@@ 5. Conclusion
As far as well-posedness of Equation 1 is concerned, we succeeded by means of a semigroup approach utilizing the theory of boundary triples; specifically, by constructing a boundary triple for the wave equation in the Lagrangian representation. This framework allows us to formulate boundary conditions that involve the displacement, its velocity and its normal stress. In particular we showed that the proposed boundary conditions induce a maximally dissipative operator. Well-posedness of Equation 1 is then a simple consequence of the Lumer–Philips theorem (and simple perturbation theory).

### v2

@@ 4. Stability
  • There exists $\Gamma\subseteq\partial\Omega\setminus\overline{\Gamma_{0}}$ open and non-empty such that $k_{2}>0$ on $\Gamma$ (i.e., there is an open set where the wave equation is damped).
@@
Making use of the results of the previous section, we have to show stability properties of the semigroup generated by $A_{\Theta}$, where $A_{\Theta}$ is the operator from the previous section, see (10), i.e., $\Theta=k_{2}$. The notion of stabiliy we have in mind is the following one:
@@
Let $H$ be a Hilbert space and let $\{T(t)\}_{t\geq 0}\subseteq\mathcal{L}(H)$ be a strongly continuous semigroup with generator $G$. The semigroup is called semi-uniformly stable if there exists a continuous non-increasing function $)f\colon[0,\infty\rparen\to[0,\infty\rparen$ satisfying $\lim_{t\to\infty}f(t)=0$, such that for every $x\in\operatorname{dom}G$:
$\lim_{t\to\infty}\lVert T(t)x\rVert_{H}\leq f(t)\lVert x\rVert_{\operatorname{dom}G}.$
@@
The notion of semi-uniform stability is nested in between strong stability, where
@@
for all $x\in H$ is demanded, and uniform stability, where
@@
We rely on the following criterium from [4, Thm. 1.1] [^14] The decay rate can be made explicit, cf. [8, Thm. 3.4] and its proof for details..
@@
Let $(T(t))_{t\geq 0}$ be a bounded, strongly continuous semigroup with generator $G$ satisfying $\upsigma(G)\cap\mathrm{i}\mathbb{R}=\emptyset$. Then $(T(t))_{t\geq 0}$ is semi-uniformly stable.
@@
The embedding of $\operatorname{dom}A_{\Theta}$ into $X$ is compact, i.e., $\operatorname{dom}A_{\Theta}\overset{\mathsf{cpt}}{\hookrightarrow}X$.
@@
$\left\lVert\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{\operatorname{dom}A}^{2}\coloneq\left\lVert\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{X}^{2}+\left\lVert A\begin{pmatrix}x_{n}\\ y_{n}\end{pmatrix}\right\rVert_{X}^{2}\leq C.$
@@
$\big\lVert\rho^{-\frac{1}{2}}y_{n}\big\rVert_{\mathrm{L}^{2}(\Omega)}^{2}+\big\lVert T^{\frac{1}{2}}\nabla x_{n}\big\rVert_{\mathrm{L}^{2}(\Omega)}^{2}+\big\lVert k_{1}^{\frac{1}{2}}\gamma_{0}x_{n}\big\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}^{2}\\ +\big\lVert\rho^{-\frac{1}{2}}\operatorname{div}T\nabla x_{n}\big\rVert^{2}_{\mathrm{L}^{2}(\Omega)}+\big\lVert T^{\frac{1}{2}}\nabla\tfrac{1}{\rho}y_{n}\big\rVert^{2}_{\mathrm{L}^{2}(\Omega)}+\big\lVert k_{1}^{\frac{1}{2}}\gamma_{0}\tfrac{1}{\rho}y_{n}\big\rVert^{2}_{\mathrm{L}^{2}(\tilde{\Gamma})}\leq C\text{.}$
@@
  • Since $(\tfrac{1}{\rho}y_{n})_{n\in\mathbb{N}}$ is bounded in $\mathrm{H}^{1}(\Omega)$, there exists a subsequence $(\tfrac{1}{\rho}y_{n(k)})_{k\in\mathbb{N}}$ that converges in $\mathrm{L}^{2}(\Omega)$ appealing to the Rellich–Kondrachov theorem ([19, Thm. 12.30]). Hence, $(\rho\tfrac{1}{\rho}y_{n(k)})_{k\in\mathbb{N}}$ converges in $\mathrm{L}^{2}(\Omega)$ as well.
  • Appealing to Rellich–Kondrachov again, boundedness of $(x_{n})_{n\in\mathbb{N}}$ in $\mathrm{H}^{1}(\Omega)$ implies existence of a subsequence that converges to $x$ in $\mathrm{L}^{2}(\Omega)$. W.l.o.g. we pass to that subsequence. To avoid the introduction of several (irrelevant) multiplicative constants, we use the symbol $\lesssim$, which stands for inequality up to a multiplicative constant independent of $n$. Note that appealing to the boundary condition prescribed by $\Theta$, we have
    $\lVert\gamma_{\nu}T\nabla x_{n}\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}=\lVert k_{1}\gamma_{0}x_{n}+k_{2}\gamma_{0}\tfrac{1}{\rho}y_{n}\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}\lesssim 1\text{.}$
    Integration by parts and the Cauchy–Schwarz inequality allow us to estimate
@@
    $\stackrel{{\scriptstyle\mathllap{\text{Int.\ by parts}}}}{{=}}\big\langle-\operatorname{div}T\nabla(x_{n}-x_{m}),x_{n}-x_{m}\big\rangle+\langle\gamma_{\nu}T\nabla(x_{n}-x_{m}),\gamma_{0}(x_{n}-x_{m})\rangle$
    $\stackrel{{\scriptstyle\mathllap{\text{C.--S.\ ineq.}}}}{{\lesssim}}\lVert x_{n}-x_{m}\rVert_{\mathrm{L}^{2}(\Omega)}+\lVert\gamma_{0}(x_{n}-x_{m})\rVert_{\mathrm{L}^{2}(\tilde{\Gamma})}$.
[…]
@@
The right-hand side is of the form $1+K$, where $K$ is a compact operator. It is a consequence of the theorem of Riesz–Schauder, cf. [30, Thm. 6.2.1], that an operator $1+K$ is injective if and only if it is surjective. Now note that the left-hand side is injective, since $\lambda-A_{\Theta}$ is injective and the resolvent is bijective. Thus $\lambda-A_{\Theta}$ is surjective. The open mapping theorem assures that $\lambda-A_{\Theta}$ is boundedly invertible. ∎
@@
If $\lambda\in\mathrm{i}\mathbb{R}$ is an eigenvalue of $A_{\Theta}$, the corresponding eigenvectors $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$ satisfy
[…]
@@
We can calculate for $\lambda\in\upsigma(A_{\Theta})$ and a corresponding eigenvector $x\in\operatorname{dom}A_{\Theta}$:
@@
$=\tfrac{1}{2}\bigl(\langle Ax,x\rangle+\langle x,Ax\rangle\bigr)-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\tfrac{1}{2}\bigl(\langle B_{1}x,B_{2}x\rangle_{\mp\frac{1}{2}}+\langle B_{2}x,B_{1}x\rangle_{\pm\frac{1}{2}}\bigr)-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\operatorname{Re}\langle B_{1}x,B_{2}x\rangle_{\mp\frac{1}{2}}-\operatorname{Re}\lambda\lVert x\rVert^{2}$
$=\operatorname{Re}\langle\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp\frac{1}{2}}-\operatorname{Re}\lambda\lVert x\rVert^{2}.$
If $\lambda\in\mathrm{i}\mathbb{R}$, we obtain
$0=\operatorname{Re}\langle\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp\frac{1}{2}}.$
Appealing to the boundary condition
$B_{1}x+\Theta B_{2}x=0\quad\Longleftrightarrow\quad k_{1}\gamma_{0}x_{1}+\gamma_{\nu}T\nabla x_{1}+k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2}=0$
and the fact that the duality can be written as an inner product if the arguments are in $\mathrm{L}^{2}(\partial\Omega)$, we infer
$0=\operatorname{Re}\langle-k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle_{\mp\frac{1}{2}}=-\operatorname{Re}\langle k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2},\gamma_{0}\tfrac{1}{\rho}x_{2}\rangle=-\lVert k_{2}^{\frac{1}{2}}\gamma_{0}\tfrac{1}{\rho}x_{2}\rVert^{2}\text{,}$
which implies $k_{2}\gamma_{0}\tfrac{1}{\rho}x_{2}=0$. Employing the boundary condition provided by $\Theta$ again, we deduce $\gamma_{\nu}T\nabla x_{1}+k_{1}\gamma_{0}x_{1}=0$. ∎
Note that the previous result says that any eigenvector $x=\begin{pmatrix}x_{1}\\ x_{2}\end{pmatrix}$ to an eigenvalue $\mathrm{i}\lambda$ satisfies $\gamma_{0}\tfrac{1}{\rho}x_{2}=0$ on $\operatorname{supp}k_{2}$, which by assumption (A6) implies that there is a non-empty open set $\Gamma$ contained in $\operatorname{supp}k_{2}$ such that $\gamma_{0}\tfrac{1}{\rho}x_{2}=0$ on $\Gamma$.
@@
$\begin{aligned} \mathrm{i}\lambda u_{1}-\tfrac{1}{\rho}u_{2}&=0\quad\text{in}\;\mathrm{H}^{1}(\Omega),\\ \mathrm{i}\lambda u_{2}-\operatorname{div}T\nabla u_{1}&=0\quad\text{in}\;\mathrm{L}^{2}(\Omega).\end{aligned}$ (11)
@@
Let $\lambda\in\mathbb{R}\setminus\{0\}$ and let us assume that $u=\begin{pmatrix}u_{1}\\ u_{2}\end{pmatrix}$ is an eigenvector to the eigenvalue $\lambda$. Then from (11) we obtain that $u_{1},u_{2}\in\mathrm{H}^{1}(\Omega)$ and plugging the second equation into the first, we obtain that $u_{1}\in\mathrm{H}^{1}(\Omega)$ has to satisfy
$\lambda^{2}\rho u_{1}+\operatorname{div}T\nabla u_{1}=0.$
Combining the first equation of (11) and Lemma 4.6 yields
@@
Given these boundary conditions, the unique continuation principle [^15] It is here that our stronger assumptions from the beginning of the section come into play. They are not required anywhere else!, cf. Appendix C, implies that $u_{1}=0$ is the only solution, which in turn implies $u_{2}=0$ and therefore $u=0$, which is a contradiction. ∎
@@
Since $\operatorname{dom}A_{\Theta}\overset{\mathrm{cpt}}{\hookrightarrow}X$ by Proposition 4.4, it suffices to show that 0 is not an eigenvalue. Suppose 0 is an eigenvalue and $x\in\operatorname{dom}A_{\Theta}$ a corresponding eigenvector. Then from (11) we immediately see $x_{2}=0$. Moreover, by Lemma 4.6
@@
$0=\langle\operatorname{div}T\nabla x_{1},x_{1}\rangle_{\mathrm{L}^{2}(\Omega)}=-\langle T\nabla x_{1},\nabla x_{1}\rangle_{\mathrm{L}^{2}(\Omega)}+\langle\gamma_{\nu}T\nabla x_{1},\gamma_{0}x_{1}\rangle_{\mp\frac{1}{2}}.$
@@
Thus $x_{1}=0$ as well, but $x=0$ is not an eigenvector. Hence, 0 cannot be an eigenvalue. ∎
@@
By Theorem 3.7, $A_{\Theta}$ generates a contraction semigroup. Theorem 4.5 shows that $\upsigma(A_{\Theta})$ consists only of eigenvalues, Lemma 4.7 and Lemma 4.8 show that the imaginary axis is contained in the resolvent set of $A_{\Theta}$. Proposition 4.3 then implies the claim. ∎
@@
We can also obtain semi-uniform stability of the full problem (1) under suitable assumptions on the perturbation
@@ 5. Conclusion
We established well-posedness of Equation 1 by constructing a boundary triple for the wave equation in the Lagrange representation. This framework allows us to formulate boundary conditions that involve the displacement, its velocity and its normal stress. In particular, we showed that the proposed boundary conditions induce a maximally dissipative operator. Well-posedness of (1) is thus a simple consequence of the Lumer–Philips theorem (and simple perturbation theory).